# Quantum Telegraph Behavior as a Result of Continuous Weak Measurement

### v1

@@ Abstract
We show that a simple model of continuous weak measurement gives rise to the telegraph switching behavior seen in experiments with single qubits. This lends support to a continuous collapse interpretation of quantum mechanics, but can also be viewed as a model of continuous detection of a steady-state process in the incoherent limit. We show explicitly that such a system obeys the Born rule for particle counting statistics, even though no particle behavior has been invoked.
@@ I Introduction
It has been argued that in quantum mechanics one can remove the idea of photons as individual corpuscles, and treat photons simply as resonances of the underlying quantum fields. The derivation of the properties of photons (and other particles) as field resonances is well known, and supports this viewpoint [1], but some experiments seem to naturally lend themselves to an interpretation in terms of individual corpuscles; namely, experiments with “clicks” or “counts.” One such experiment is when a wave with amplitude of $N=1$ is sent to multiple photon detectors, and only one of them clicks, even if they are spacelike separated. Another is when a continuous electromagnetic field impinges on a single atom or two-level system, and that system responds with “telegraph noise;” that is, the state of the system jumps suddenly and randomly between two states, in what can be seen as discrete photon absorption and photon emission events (e.g., Refs. [2, 3]).
@@
It is common to interpret the stochastic nature of clicks and counts in quantum mechanics as arising from “measurement.” In the standard model of measurement, the clicks or counts seen in such experiments are the result of an intrinsically random, non-unitary process off the books of the equations of quantum mechanics; for example, a process instigated by human knowledge. However, two recent developments in the theory of measurement allow us to take a different perspective. The first is the large body of work on “weak” measurement, [4], which shows that what we encounter as all-or-nothing measurements can be seen as the result of a large number of weak, partial measurements. These can be represented by a differential operator [5, 6] that can be added to the Hamiltonian of the Schrödinger equation.
The second recent development is a variation of spontaneous collapse theory [5] that shows that the same results can be obtained by dismissing the measurement apparatus altogether, and simply positing the differential form of weak measurements as a universal stochastic noise term, with a noise level proportional to actual physical random fluctuations in the system.
@@
It has already been shown [6] that the clicking of one detector out of the midst of many can be interpreted as follows: an input optical plane-wave puts multiple atoms into a superposition of both the upper and lower states of a two-level system, and then a stochastic process causes one of the atoms to end up in its upper state, while jamming all of the others into the lower state; this can happen suddenly, with a measurement time that is not strongly affected by the number of atoms used as detectors. In other words, the random “click” at one atom comes ultimately not from a pre-existing photon corpuscle, but from a random walk due to quantum noise acting on atoms with discrete excitation states.
@@
In this paper we show that the same type of analysis can be applied to the second type of experiment mentioned above, namely telegraph noise. A single two-level system (qubit) is subject to a continuous, steady-state, incoherent pump, but shows stochastic jumps between its upper and lower state, with statistics that track with the expected probability according to the Born measurement rule. No photons are invoked at all to see this behavior.
As discussed above, this lends itself to the interpretation of universal spontaneous collapse without a need for detectors to be treated as special devices off the books of quantum dynamics. However, it is also a useful result in its own right, even if one sticks to standard measurement-device theory. In any qubit system, low-level continuous pumping with energy loss and continuous measurement will result in telegraph noise, jumping between its two states, which obeys the quantum Born rule.
@@ II The two-level model with continuous pumping and decay
where $\vec{U}$ is the standard Bloch vector in the rotating frame, $T_{2}$ is a decoherence time and $T_{1}$ is an intrinsic energy-loss time, $\tilde{\omega}$ is the detuning of the pump frequency from the qubit resonance frequency, and $\omega_{R}$ is the Rabi oscillation frequency, proportional to the intensity of the pump wave. In the case of an incoherent pump, a term can be added to the last equation that acts oppositely to a $T_{1}$ loss process, namely $G(1-U_{3})$, where $G$ is the incoherent pump rate, which is either zero or positive, and the $(1-U_{3})$ factor accounts for Pauli exclusion, preventing the upper state from having occupation greater than 1.
If we assume fully incoherent pumping, this gives us
@@
In Appendix A we justify this prediction by analyzing the behavior of $U_{3}$.
@@
where $\varepsilon$ is a small, fluctuating term that gives, if acting alone, a Martingale random walk of $U_{3}$. This random walk has been shown [7] to give trapping (“collapse”) at the values $U_{3}=\pm 1$ for states that begin in a superposition between these two values, with probability in agreement with the standard Born rule for quantum measurements.
If this stochastic term acts alone on a qubit, the random walk of the Bloch vector will end up at the top or the bottom of the sphere and stay there. However, if we add together the two terms (2) and (5), we will have dynamical instability. It can easily be seen that the term (5) never leads the Bloch vector to exactly $U_{3}=1$ or $U_{3}=-1$ (as in the quantum Zeno effect [4, 8, 9]; the step size of the random walk decreases as $U_{3}$ approaches one of these end points, so that it becomes quite close, and trapped, so to speak, but never quite at the end. In this context, the “force” (2) has the effect of pushing the Bloch vector away from the ends: at $U_{3}=1$, ${\partial U_{3}}/{\partial t}=-2/T$, and at $U_{3}=-1$, ${\partial U_{3}}/{\partial t}=2G$. Therefore there is a nonzero chance that the Bloch vector will escape an end point, and undergo a new random walk in the middle of the sphere. As shown in Appendix A, when this occurs, the statistical prediction (4) can justified mathematically.
This model is essentially the same as that considered in the context of weak measurement theory with coherent-wave pumping (see, e.g., Ref. [4], Chapters 5 and 6, and references therein), but with an incoherent pump. By casting this model only in terms of the non-unitary term (5), we show that weak measurement theory is not neeeded, however, only universal continuous collapse, which converges to the predictions of weak measurement theory in all known cases. In particular, the term (5) is all that is sufficient for particle-counting clicks that follow the Born rule of counting statistics.
@@ IV.1 Weak measurement
Telegraph behavior has been studied in the weak measurement literature, as well as the quantum Zeno effect [4, 8, 9]. A potential source for switching behavior is from the randomness introduced to both the decoherence and incoherence by the measurement process, via the following set of equations [4]:
@@ V Dependence of the Long-Term Behavior of $U_{3}$ on $\alpha$
Figures 1 and 3 show that depending on the strength of the quantum noise parameter $\alpha$, there can be three qualitatively different behaviors. Large $\alpha$ leads to telegraph behavior, intermediate values of $\alpha$ cause noisy trajectories, while small values of $\alpha$ cause the process to behave mostly deterministically. Since we can only distinguish these behaviors in the long run (after $t\approx GT$, see Appendix A), what is important is the behavior of $U_{3}$ in the $t\to\infty$ limit. Mathematically, that is what the stationary probability density $\rho(y)$ describes. This density would satisfy $\mathbb{P}(a\leqslant U_{3}\leqslant b)=\int_{a}^{b}\rho(y)\,dy$ for all large times $t$.
We used the Fokker-Planck equation (see Ref. [17], Section 4.4) to derive the density $\rho(y)$ for our process $U_{3}$. (More details can be found in Appendix C.) We found that
@@ VI Conclusions
Appendix A A Justification of the Statistics

### v2

# Quantum Telegraph Behavior Without Photons

@@ Abstract
We show that a simple model of non-Hermitian noise gives rise to the telegraph switching behavior seen in experiments with single qubits, without any reference to the existence of photons as corpuscles. This lends support to a continuous collapse interpretation of quantum mechanics, but can also be viewed as a model of continuous detection of a steady-state process in the incoherent limit. We show explicitly that such a system obeys the Born rule for particle counting statistics, even though no particle behavior has been invoked at any point in the calculation.
@@ I Introduction
It has been argued that in quantum mechanics one can remove the idea of photons as individual corpuscles, and treat photons simply as resonances of the underlying quantum fields. The derivation of the properties of photons (and other particles) as field resonances is well known, and supports this viewpoint (for a review, see Ref. 1), but some experiments seem to naturally lend themselves to an interpretation in terms of individual corpuscles; namely, experiments with “clicks” or “counts.” One such experiment is when a wave with amplitude of $N=1$ is sent to multiple photon detectors, and only one of them clicks, even if they are spacelike separated. Another is when a continuous electromagnetic field impinges on a single atom or two-level system, and that system responds with “telegraph noise;” that is, the state of the system jumps suddenly and randomly between two states, in what can be seen as discrete photon absorption and emission events. Telegraph noise has been observed in trapped ions [2, 3, 4], cavity fields [5], molecules [6], and superconducting artificial atoms [7], just to list a few.
@@
It is common to interpret the stochastic nature of clicks and counts in quantum mechanics as arising from measurement of single photons. In the standard model of measurement, the clicks or counts seen in such experiments are the result of an intrinsically random, non-unitary process, which cannot be accounted for by the Schrödinger equation or by any equations derived from it, such as Lindbladian dynamics. For example, in the Copenhagen interpretation, such a non-unitary process is instigated by human knowledge. However, two recent developments in the theory of measurement allow us to take a different perspective. The first is the large body of work on “weak” measurement (for a review, see, e.g., Ref. 8), which shows that what we encounter as all-or-nothing measurements can be seen as the limiting behavior of a large number of weak, partial measurements. As shown in prior work [9, 10], the effect of weak measurements can be represented by a differential operator that can be added to the Hamiltonian of the Schrödinger equation.
A second body of work that has arisen in recent decades is spontaneous collapse theory, which has had several variants such as continuous spontaneous localization (CSL) [11, 12, 13, 14, 15, 16], the Ghirardi-Rimini-Weber (GRW) model [17, 18, 19], and nonlocal stochastic collapse [20, 21]. A recent variation [22, 23, 9, 10] takes a similar approach, but argues that no net violation of energy conservation is needed, if a simple non-Hermitian term is added to the Schrödinger equation. It has been shown [9, 10] that the non-Hermitian different operator added to the Schrödinger equation in this approach is identical to that derived from weak measurement theory. Therefore, one can take all the mathematical results of weak measurement theory, but dismiss the measurement apparatus altogether, and simply posit this differential operator as a universal stochastic noise term, with a noise level proportional to actual physical random fluctuations in the system.
@@
It has already been shown [10] that the clicking of one detector out of the midst of many can be interpreted as follows: an input optical plane-wave puts multiple atoms into a superposition of both the upper and lower states of a two-level system, and then a stochastic process causes one of the atoms to end up in its upper state, while jamming all of the others into the lower state; this can happen suddenly, with a measurement time that is not strongly affected by the number of atoms used as detectors. In other words, the random “click” at one atom comes ultimately not from a pre-existing photon corpuscle, but from a random walk due to quantum noise acting nonlocally on many atoms with discrete excitation states; it is not necessary at all to have a photon as a corpuscle to get just one of the detectors to click.
@@
In this paper we show that the same type of analysis can be applied to the second type of experiment mentioned above, namely telegraph noise. A single two-level system (qubit) is subject to a continuous, steady-state, incoherent pump, but shows stochastic jumps between its upper and lower state, with statistics that track with the expected probability according to the Born measurement rule. No photons as corpuscles are invoked to obtain this behavior.
Prior work on telegraphing behavior in the context of weak measurement theory [8, 24, 25, 26] has assumed the interpretation that the “measurements” are, at the end of the day, gaining human knowledge about what photon corpuscles are doing. However, the entire effect of the “measurements” is to introduce stochastic noise, which, as we show here, reproduces the quantum Born rule for particle counting, namely, that the probability of a count is proportional to the square of the amplitude of the wave function at the detector. In any two-state system, low-level continuous pumping with energy loss and noise will result in telegraph noise, jumping between its two states, which obeys the Born rule. Therefore it is appealing to simply posit the quantum noise term as fundamental.
As discussed in Section IV, the quantum mechanical results obtained in this case are similar to classical telegraphing systems, but have important differences.
@@ II The two-level model with continuous pumping and decay
where $\vec{U}$ is the standard Bloch vector in the rotating frame, $T_{2}$ is a decoherence time and $T_{1}$ is an intrinsic energy-loss time, $\tilde{\omega}$ is the detuning of the pump frequency from the qubit resonance frequency, and $\omega_{R}$ is the Rabi oscillation frequency, proportional to the intensity of the pump wave.
In the case of an incoherent pump, a term can be added to the last equation (see Appendix A) that acts oppositely to a $T_{1}$ loss process, namely $G(1-U_{3})$, where $G$ is the incoherent pump rate, which is either zero or positive, and the $(1-U_{3})$ factor accounts for Pauli exclusion, preventing the upper state from having occupation greater than 1. If we assume fully incoherent pumping, this gives us
@@
In Appendix B we justify this prediction by analyzing the behavior of $U_{3}$.
@@
where $\varepsilon$ is a small, fluctuating term that gives, if acting alone, a Martingale random walk of $U_{3}$. As discussed in the Introduction, this term can either be justified as a fundamental postulate of spontaneous collapse theory, or as the differential limiting behavior of continuous weak measurement. The random walk generated by this term has been shown [23] to give trapping (“collapse”) at the values $U_{3}=\pm 1$ for states that begin in a superposition between these two values, with probability in agreement with the standard Born rule for quantum measurements.
If this stochastic term acts alone on a qubit, the random walk of the Bloch vector will end up at the top or the bottom of the sphere and stay there. However, if we add together the two terms (2) and (5), we will have dynamical instability. It can easily be seen that the term (5) never leads the Bloch vector to exactly $U_{3}=1$ or $U_{3}=-1$ (as in the quantum Zeno effect [8, 25, 24]; the step size of the random walk decreases as $U_{3}$ approaches one of these end points, so that it becomes quite close, and trapped, so to speak, but never quite at the end. In this context, the “force” (2) has the effect of pushing the Bloch vector away from the ends: at $U_{3}=1$, ${\partial U_{3}}/{\partial t}=-2/T$, and at $U_{3}=-1$, ${\partial U_{3}}/{\partial t}=2G$. Therefore there is a nonzero chance that the Bloch vector will escape an end point, and undergo a new random walk in the middle of the sphere. As shown in Appendix B, when this occurs, the statistical prediction (4) can justified mathematically.
This model is essentially the same as that considered in the context of weak measurement theory with coherent-wave pumping (see, e.g., Ref. [8], Chapters 5 and 6, and references therein), but with an incoherent pump. By casting this model only in terms of the non-unitary term (5), we show that the full apparatus of weak measurement theory is not neeeded, however; simply positing a universal noise term of the form used here converges to the predictions of weak measurement theory in all known cases. That is, the term (5) is by itself sufficient for particle-counting clicks that follow the Born rule of counting statistics.
@@ IV.1 Weak measurement
Telegraph behavior has been studied in the weak measurement literature, as well as the quantum Zeno effect [8, 25, 24, 29]. A potential source for switching behavior is from the randomness introduced to both the decoherence and incoherence by the measurement process, via the following set of equations [8]:
@@ V Dependence of the Long-Term Behavior of $U_{3}$ on $\alpha$
Figures 1 and 3 show that depending on the strength of the quantum noise parameter $\alpha$, there can be three qualitatively different behaviors. Large $\alpha$ leads to telegraph behavior, intermediate values of $\alpha$ cause noisy trajectories, while small values of $\alpha$ cause the process to behave mostly deterministically. Since we can only distinguish these behaviors in the long run (after $t\approx GT$, see Appendix B), what is important is the behavior of $U_{3}$ in the $t\to\infty$ limit. Mathematically, that is what the stationary probability density $\rho(y)$ describes. This density would satisfy $\mathbb{P}(a\leqslant U_{3}\leqslant b)=\int_{a}^{b}\rho(y)\,dy$ for all large times $t$.
We used the Fokker-Planck equation (see Ref. [35], Section 4.4) to derive the density $\rho(y)$ for our process $U_{3}$. (More details can be found in Appendix D.) We found that
@@ VI Conclusions
Appendix A Derivation of the driven-dissipative Bloch equations
We start with a generic two-level system with state $|\psi\rangle=\alpha|g\rangle+\beta|e\rangle$. The density matrix for this system is
(8) $\hat{\rho}=\left(\begin{array}[]{cc}|\alpha|^{2}&\alpha^{*}\beta\\ \beta^{*}\alpha&|\beta|^{2}\end{array}\right),$
In the standard perturbation approach [1], we write the Hamiltonian $H=H_{0}+\hat{V}$, with the first-order evolution given by
(9) $\frac{d}{dt}\langle\hat{\rho}_{i,j}\rangle$ $=$ $\frac{i}{\hbar}\displaystyle\sum_{n}|\alpha_{n}|^{2}\langle[H_{0}+\hat{V},\hat{\rho}_{i,j}]\rangle$
and the second-order evolution given by the Lindbladian,
[…]
where the final $\delta$-function gives the energy conservation for any initial and final state.
The Hamilton for the two-state system coupled to a driving field is
(13) $H_{0}=\left(\begin{array}[]{cc}\hbar\omega_{0}&\displaystyle\frac{\hbar\omega_{R}}{2}e^{-i\omega t}\\ \displaystyle\frac{\hbar\omega_{R}}{2}e^{i\omega t}&0\end{array}\right),$
while the coupling to the environment is given by
(14) $\hat{V}=\sum_{q}A_{q}(a_{q}b^{\dagger}_{e}b_{g}+a_{q}^{\dagger}b^{\dagger}_{g}b_{e}),$
where $b^{\dagger}_{i}$, $b_{i}$ are the fermionic creation and destruction operators for state $i$, $a^{\dagger}_{q}$, $a_{q}$ are bosonic creation and destruction operators for a photon in state $q$, and $A_{q}$ is a constant giving the strength of the interaction to emit or absorb a photon. This term vanishes in the first-order evolution term (9), since no phase coherence of the photons in the environment is assumed, apart from the driving laser, already accounted for in (13).
For spontaneous emission, we assume no photons initially. Writing $\hat{\rho}_{ij}=\langle b^{\dagger}_{i}b_{j}\rangle$, and using $N_{e}+N_{g}=1$ and $(b_{i})^{2}=(b^{\dagger}_{i})^{2}=0$, we obtain
$\langle 2(a_{q}b^{\dagger}_{e}b_{g})b^{\dagger}_{e}b_{e}(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})-(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})b^{\dagger}_{e}b_{e}-b^{\dagger}_{e}b_{e}(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})\rangle=-2\langle N_{e}\rangle$
$\langle 2(a_{q}b^{\dagger}_{e}b_{g})b^{\dagger}_{g}b_{g}(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})-(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})b^{\dagger}_{g}b_{g}-b^{\dagger}_{g}b_{g}(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})\rangle=2\langle N_{e}\rangle$
$\langle 2(a_{q}b^{\dagger}_{e}b_{g})b^{\dagger}_{e}b_{g}(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})-(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})b^{\dagger}_{e}b_{g}-b^{\dagger}_{e}b_{g}(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})\rangle=-\langle b_{e}^{\dagger}b_{g}\rangle$
$\langle 2(a_{q}b^{\dagger}_{e}b_{g})b^{\dagger}_{g}b_{e}(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})-(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})b^{\dagger}_{g}b_{e}-b^{\dagger}_{g}b_{e}(a_{q}b_{e}^{\dagger}b_{g})(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})\rangle=-\langle b_{g}^{\dagger}b_{e}\rangle.$
Defining
$U_{1}$ $=$ $\beta^{*}\alpha+\alpha^{*}\beta$
$U_{2}$ $=$ $i(\beta^{*}\alpha-\alpha^{*}\beta)$
(16) $U_{3}$ $=$ $|\beta|^{2}-|\alpha|^{2},$
and $|\alpha|^{2}+|\beta|^{2}=1$, we then have for the Lindbladian term
(17) $\frac{\partial U_{3}}{\partial t}$ $=$ $-\frac{1}{T_{1}}2|\beta|^{2}=-\frac{1}{T_{1}}(U_{3}+1),$
where $T$ is the characteristic photon emission time, and
$\frac{\partial U_{1}}{\partial t}$ $=$ $-\frac{1}{2T_{1}}U_{1},$
(18) $\frac{\partial U_{2}}{\partial t}$ $=$ $-\frac{1}{2T_{1}}U_{2}.$
These three equations give the standard dissipation terms in the Bloch equations.
We can now adopt the rotating frame and writing $|\psi\rangle=\alpha^{\prime}|v\rangle+\beta^{\prime}e^{-i\omega_{0}t}|c\rangle$. Writing the Bloch vector in terms of $\alpha^{\prime}$ and $\beta^{\prime}$, for the case when the coherent driving field frequency $\omega$ equals $\omega_{0}$, Schrödinger’s equation for the Hamiltonian (13) plus the dissipative terms give the standard equations in the rotating frame,
$\displaystyle\frac{\partial U^{\prime}_{1}}{\partial t}$ $=$ $\displaystyle-\frac{U^{\prime}_{1}}{T_{2}}$
$\displaystyle\frac{\partial U^{\prime}_{2}}{\partial t}$ $=$ $\displaystyle-\frac{U^{\prime}_{2}}{T_{2}}-\omega_{R}U^{\prime}_{3}$
(19) $\displaystyle\frac{\partial U^{\prime}_{3}}{\partial t}$ $=$ $\displaystyle-\frac{U^{\prime}_{3}+1}{T_{1}}+\omega_{R}U^{\prime}_{2},$
with $T_{2}=2T_{1}$. This assumes that the the photon emission gives both relaxation from the upper state and also decoherence; it is also possible to have additional decoherence processes that do not lead to relaxation, giving an effective $T_{2}$ that is shorter.
We can now take exactly the same approach to account for incoherent pumping. In this case, analogous to (15), we start with an initial state with one photon in state $q$, and write
$\langle 2(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})b^{\dagger}_{e}b_{e}(a_{q}b^{\dagger}_{e}b_{g})-(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})b^{\dagger}_{e}b_{e}-b^{\dagger}_{e}b_{e}(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})\rangle=2\langle N_{g}\rangle$
$\langle 2(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})b^{\dagger}_{g}b_{g}(a_{q}b^{\dagger}_{e}b_{g})-(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})b^{\dagger}_{g}b_{g}-b^{\dagger}_{g}b_{g}(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})\rangle=-2\langle N_{g}\rangle.$
These give us, for the same definitions as above,
(21) $\frac{\partial U_{3}}{\partial t}$ $=$ $2G|\alpha|^{2}=G(1-U_{3}),$
which is the incoherent pumping term used in the main text. Similarly, the off-diagonal terms are
$\langle 2(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})b^{\dagger}_{e}b_{g}(a_{q}b^{\dagger}_{e}b_{g})-(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})b^{\dagger}_{e}b_{g}-b^{\dagger}_{e}b_{g}(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})\rangle=-\langle b_{e}^{\dagger}b_{g}\rangle,$
$\langle 2(a^{\dagger}_{q}b^{\dagger}_{g}b_{e})b^{\dagger}_{g}b_{e}(a_{q}b^{\dagger}_{e}b_{g})-(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})b^{\dagger}_{g}b_{e}-b^{\dagger}_{g}b_{e}(a^{\dagger}_{q}b_{g}^{\dagger}b_{e})(a_{q}b_{e}^{\dagger}b_{g})\rangle=-\langle b_{g}^{\dagger}b_{e}\rangle$
which give decoherence due to the incoherent pumping. This can then be added to the first two equations of (1) to give an adjusted $T_{2}$ time.
Note that the Lindbladian approach used here gives irreversible behavior, namely deterministic decay, but cannot give stochastic behavior, that is, different results for the same initial conditions. This is fundamentally because any approach derived from unitary quantum mechanics, including the Lindbladian term used here, acts only on the many-body wave function of the system plus its enviroment, and therefore gives deterministic behavior. This is fundamentally the same result as found in Ref. 37.
Appendix B Justification of the Statistics